what is the length of the base if the height is 5 less than the base and the area of the triangle is 21?
Answer by nyc_function(2741)

A = 1/2 times base times height
Let A = 21
Let h = height = (x - 5)
Let b = base = x
21 = 1/2 times x times (x - 5)
I gave you the equation. Can you finish now?